0.5 kg of an ideal gas undergoes an isentropic process from 657 kPa gage and a volume of 0.017 cubic meter to a final volume of 0.1 cubic meter. If Cv = 0.384 kJ/kg-K and Cp = 0.52 kJ/kg-K, calculate:

a) the final temperature in deg. C

b) the final pressure in kPa

c) the change in enthalpy in kJ

d) the work nonflow in kJ


1
Expert's answer
2021-12-02T10:08:04-0500

a)

"\nu=\frac{C_p-C_V}R=16.4~mol,"

"\gamma=1+\frac{\nu R}{C_V}=1.36,"

"T_1=\frac{p_1V_1}{\nu R}=82~K,"

"T_2=T_1(\frac{V_1}{V_2})^{\gamma-1}=43~K=-230°C,"

b)

"p_2=\frac{\nu RT_2}{V_2}=58.6~kPa,"

c)

"\Delta H=mC_p\Delta T=10.2~kJ,"

d)

"A=\nu C_V\Delta T=246~kPa."